Ted cleaned the first floor in 4 hours. With Alex helping, they did the second floor in 3 hours. How long would it take Alex to do both floors alone?
Answer by stanbon(75887)

Ted cleaned the first floor in 4 hours. With Alex helping, they did the second floor in 3 hours. How long would it take Alex to do both floors alone?
--------
Ted rate: 1/4 job/hr
Together rate: 1/3 job/hr
Alex rate: 1/x job/hr
--------------------------
Equation:
rate + rate = together rate
1/4 + 1/x = 1/3
------
Multiply thru by 12x to get:
3x + 12 = 4x
x = 12 (time it would take Alex to do the job alone)
====================
Cheers,
Stan H.